OpenStudy (anonymous):

Question 5 (Multiple Choice Worth 2 points) (03.08) Solve -2x2 +3x - 9 = 0. x equals quantity of 3 plus or minus 3i square root of 7 all over 4 x equals quantity of 3 plus or minus 9i square root of 7 all over 4 x equals quantity of negative 3 plus or minus 3i square root of 7 all over 4 x equals quantity of negative 3 plus or minus 9i square root of 7 all over 4

OpenStudy (anonymous):

Im new here but i really need help :(

OpenStudy (sleepyjess):

Hey there! Welcome to OpenStudy!

OpenStudy (anonymous):

thanks

OpenStudy (sleepyjess):

Just wondering, is this a quiz?

OpenStudy (anonymous):

sort of yea

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

Is it like a lesson review? I have those :)

OpenStudy (anonymous):

Its one of those quizes on flvs that lets you redo the assignment 3 times

OpenStudy (sleepyjess):

Yeah, it is pretty much a lesson review

OpenStudy (dtan5457):

You will want to use the quadratic formula a=-2 b=3 c=-9

OpenStudy (dtan5457):

\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (dtan5457):

just plug them in and simplify to get x

OpenStudy (anonymous):

okay thanx ill try it out then come back to you @dtan5457

OpenStudy (anonymous):

so inside the radical is it -63? @dtan5457

OpenStudy (dtan5457):

yes

OpenStudy (anonymous):

yayy so uhh idk how to simplify the rest.

OpenStudy (dtan5457):

-2x2 +3x - 9 = 0 are you sure there is a negative in front of that two

OpenStudy (anonymous):

ik its a complex solution but thats about it @dtan5457

OpenStudy (dtan5457):

one of your answer choices has to be over -4, not positive 4

OpenStudy (anonymous):

Thats basically all the answer choices they gave me

OpenStudy (dtan5457):

oh wait never mind, i see it.

OpenStudy (dtan5457):

\[\frac{ -3\pm \sqrt{-63} }{ -4 }\]

OpenStudy (dtan5457):

when you plug it all in you get that, which of those simplify to that?

OpenStudy (anonymous):

Yea thats what i have,and wdym

OpenStudy (dtan5457):

you can further simplify that radical of -63

OpenStudy (anonymous):

dont you get a decimal?

OpenStudy (dtan5457):

what perfect squares multiplied with any number equals 63? 4,9,16,25,36, etc

OpenStudy (dtan5457):

you don't actually find the square root of -63

OpenStudy (dtan5457):

but it can be put into simplified form

OpenStudy (dtan5457):

how about i give you an example

OpenStudy (anonymous):

No thank you lol

OpenStudy (dtan5457):

lol, do you know what i'm talking about though? its extremely important you get the concept..

OpenStudy (acxbox22):

there is no solution if the number under the radical is negative in the quadratic formula

OpenStudy (dtan5457):